I am working on substitution an elimination.I can figure out how to do the problems when they are in the right form but I don't know how to solve it when the problem is set up different. example... 2x-y=7 -7y=2x-31 On my paper the two equations are lined up over each other but I can't get it like that on here.
Answer by TimothyLamb(4379)

2x - y = 7
-7y = 2x - 31
---
2x - y = 7
-2x - 7y = -31
---
add
---
-8y = -24
---
y = 3
2x - y = 7
2x - 3 = 7
2x = 10
x = 5
---
Answer:
x = 5
y = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php